### v1

# Watermarking Language Models through Language Models

Abstract
This paper presents a novel framework for watermarking language models through prompts generated by language models. The proposed approach utilizes a multi-model setup, incorporating a Prompting language model to generate watermarking instructions, a Marking language model to embed watermarks within generated content, and a Detecting language model to verify the presence of these watermarks. Experiments are conducted using ChatGPT and Mistral as the Prompting and Marking language models, with detection accuracy evaluated using a pretrained classifier model. Results demonstrate that the proposed framework achieves high classification accuracy across various configurations, with 95% accuracy for ChatGPT, 88.79% for Mistral. These findings validate the and adaptability of the proposed watermarking strategy across different language model architectures. Hence the proposed framework holds promise for applications in content attribution, copyright protection, and model authentication.
Keywords Watermarking  $\cdot$ Large Language Model
1 Introduction
Language model watermarking refers to the practice of embedding identifiable markers or signatures within the outputs of language models. These watermarks are designed to be imperceptible to end-users but detectable by an authorized system, allowing the origin of the content to be traced back to a particular model or model owner. This process typically involves embedding patterns or codes within the generated text in such a way that they do not affect its quality or meaning while enabling traceability and authentication. Recent advancements in deep learning have opened new opportunities for incorporating these watermarks efficiently within large language models (LLMs).
The need for language model watermarking arises due to the growing prevalence of generative AI systems in real-world applications, where issues such as content ownership, intellectual property protection, and accountability are becoming increasingly critical. As language models are deployed in diverse settings, from automated content generation to conversational agents, ensuring the integrity of the generated text is vital. Watermarking can help identify misuse of models, detect AI-generated content, and enforce responsible usage, particularly when it comes to protecting copyrighted material, preventing deepfake text, or monitoring compliance with ethical guidelines.
Recent advancements in watermarking techniques for LLMs have focused on two primary approaches: embedding watermarks directly into generated text and embedding watermarks within the LLM architecture itself. Text-based watermarking methods typically alter token probabilities or manipulate token selections during generation to make watermarks detectable by specific algorithms while maintaining imperceptibility for users. Notable contributions include signal embedding through token manipulation [1, 2], cryptographically inspired undetectable watermarks [3], and adaptive strategies based on token entropy [4]. While these methods achieve various levels of robustness, detectability, and minimal impact on text quality, model-based watermarking techniques embed watermarks within the internal structure of LLMs [5, 6], protecting models against tampering and unauthorized fine-tuning. These methods ensure the integrity of model weights, prevent unauthorized use, and safeguard intellectual property. Despite the progress, challenges remain in achieving dynamic, robust, and imperceptible watermarking that can adapt to different text contexts.
In this work, we propose a novel dynamic watermarking method that leverages the internal capabilities of language models to generate and insert watermarks via prompts during content generation. Unlike existing methods that statically embed watermarks into text or model parameters, our approach dynamically adjusts the watermarking based on the model’s output, ensuring flexibility and robustness across various applications. Our advantages include: (1) We introduce a language model-guided watermarking mechanism that dynamically adapts to the generation process, and (2) We demonstrate that our approach maintains high text quality and detection accuracy without requiring access to the model’s parameters or API.
2 Related Work
Watermarking techniques for LLMs have gained traction as essential tools for safeguarding intellectual property and authenticating AI-generated content. Broadly, these methods can be categorized [7] into two types: watermarking embedded in the generated text and watermarking embedded directly in the LLM’s architecture.
Watermarking in Generated Text. Watermarking in generated text has been an effective approach to identify AI-generated content while protecting intellectual property. Several methods focus on embedding signals directly into the generated outputs of LLMs, making these watermarks detectable by algorithms but imperceptible to humans.
Kirchenbauer et al. [1] introduced a watermarking framework for large language models that embeds signals into generated text, which remain invisible to humans but are algorithmically detectable from a short span of tokens. Their method leverages a randomized selection of "green" tokens to softly promote their use during sampling, thus embedding the watermark with negligible impact on text quality. The watermark detection relies on a statistical test with interpretable p-values, without requiring access to the model’s API or parameters. Zhao et al. [2] proposed the Unigram-Watermark, a method for watermarking LLM-generated text by grouping tokens in a fixed manner. They provide a theoretical framework assessing watermark effectiveness, robustness, and tolerance to text editing and paraphrasing. Experiments demonstrate detection accuracy with small impact on text quality. Wang et al. [8] introduced Codable Text Watermarking for LLMs, enabling watermarks to carry multi-bit customizable information, such as model version or user ID. They use a token-based Balance-Marking method, which employs a proxy language model to split the vocabulary into balanced parts. Their evaluation system includes metrics such as watermark success rate, robustness, and impact on text quality. Christ et al. [3] proposed a cryptographically-inspired method for embedding undetectable watermarks in language model outputs by subtly altering token probabilities. Their approach requires a secret key for detection and ensures no observable change to the output distribution, maintaining the quality of the generated text. The method is based on the existence of one-way functions and is designed to remain undetectable even under adaptive querying. Kuditipudi et al. [9] proposed a watermarking methodology for autoregressive language models that maps tokens to a sequence of random numbers generated with a randomized key. Watermarked text is generated by aligning the output tokens to this random number sequence. The method was applied to multiple models, demonstrating robustness to paraphrasing attacks and reliable detection from a small number of tokens, though detection performance varied across models. Guan et al. [10] introduced CodeIP, a multi-bit watermarking technique for code generated by LLMs that inserts watermarks by manipulating tokens to preserve provenance details such as vendor ID. The approach ensures syntactical correctness by constraining the token prediction process with a type predictor. Experiments across multiple programming languages demonstrated the method’s effectiveness in watermarking while maintaining code correctness. Liu and Bu [4] proposed an adaptive watermarking strategy for LLM-generated text. Their method selectively adds watermarks to high-entropy token distributions, while leaving low-entropy tokens untouched. They adaptively scale the output logits based on semantic embeddings of previously generated text, avoiding the use of fixed token lists. Experiments demonstrate robustness and security, with text quality comparable to un-watermarked models. In contrast to token probability manipulation, Munyer et al. [11] developed DeepTextMark, which embeds watermarks by substituting lexical synonyms within the generated text. This method relies on subtle lexical changes, such as synonym replacements, that preserve the text’s meaning while embedding a watermark, ensuring both robustness and imperceptibility. These diverse approaches illustrate a growing focus on balancing robustness, subtlety, and quality in watermarking AI-generated text.
Watermarking within Language Model Architectures. An emerging area of research focuses on embedding watermarks directly within the parameters of LLMs to safeguard the integrity of the models themselves. Unlike text-based watermarks, these techniques target the internal structure of LLMs, ensuring protection against tampering and misuse at the model level.
Li et al. [5] proposed a watermarking strategy that embeds watermarks in the quantization process of LLMs. Their approach allows the watermark to be active in fp32 mode while remaining hidden in int8 mode, preventing further fine-tuning of the model. The method was successfully applied to open-source models such as GPT-Neo and LLaMA, providing a potential direction for protecting model weights. Liu et al. [6] introduced TextMarker, a backdoor-based watermarking technique that protects private information embedded in training data through membership inference. The method requires marking a small subset of the data and is effective under black-box access assumptions. Experiments show that marking as little as 0.1% of the data enables effective membership inference with minimal impact on model performance. Xu et al. [12] introduced a lightweight fingerprinting technique for LLMs using instruction tuning. A private key is implanted as an instruction backdoor, causing the model to generate specific text when the key is present. The method is shown to be effective across 11 LLMs without affecting normal model behavior, while also providing robustness against fingerprint guessing and parameter-efficient training. Huang et al. [13] introduced multi-designated detector watermarking (MDDW) for LLMs, allowing specific detectors to identify watermarked outputs while maintaining output quality for regular users. The method is based on multi-designated verifier signatures (MDVS) and supports an optional claimability feature, enabling model providers to assert ownership of outputs. Their implementation demonstrates flexibility and satisfactory performance metrics. Ren et al. [14] proposed Robust and Imperceptible Watermarking (RIW) for LLM-generated text, which leverages token prior probabilities to balance detectability and imperceptibility. The method partitions tokens into two groups based on their prior probabilities and applies different embedding strategies. A ’voted z-test’ is used for watermark detection. Experimental results show that RIW improves robustness and text quality across multiple LLMs. Yoo et al. [15] proposed a Multi-bit Watermark via Position Allocation, enabling traceable information embedding in LLM outputs. The method allocates tokens across different parts of the text, allowing for longer message embedding in high-corruption settings while maintaining robustness and low latency. Their approach supports both zero-bit detection and multi-bit watermarking without requiring model access or fine-tuning, preserving text quality. These methods illustrate the ongoing development of watermarking techniques directly within LLM architectures, balancing robustness, subtlety, and performance while protecting intellectual property and preventing model misuse.
Prompt-driven Watermarking. Yao et al. [16] introduced the concept of prompt-driven watermarking for copyright protection. In this approach, a prompt is used to inject watermarks into the output of a language model. The watermark can then be verified by prompting the model again. The technique highlights the flexibility of using prompts to control and verify the presence of watermarks.
Many existing methods for watermarking in generated text and embedding watermarks within language model architectures often require access to the model’s API or internal parameters. This requirement can limit their usability in scenarios with restricted API access or proprietary models. Additionally, these methods tend to be static, meaning they apply the same watermarking strategy regardless of the specific content or context of user requests, which reduces their adaptability. In addition, existing prompt-driven method uses prompts as a mechanism to trigger or verify the watermark for copyright protection and traceability. The concept of prompting a model to watermark itself continuously or systematically across all outputs (without user intervention) is uncharted. To this end, we propose using prompts as a self-guided process initiated by the model.
3 The Proposed Method
In this section, we introduce a novel approach to watermarking language models (LMs) using a combination of language models. The proposed framework includes a prompting language model (Prompting LM), a marking language model (Marking LM), and a detecting language model (Detecting LM). The method is designed to dynamically generate, embed, and detect watermarks within text generated by LM.
Figure 1: The proposed method.
As illustrated in Figure 1, this proposed method leverages multiple LMs to implement a dynamic, adaptive watermarking technique. The prompting LM generates system instructions tailored to the user’s input, the marking LM watermarks the output based on these instructions, and the detecting LM uses binary classification to identify watermarked text.
3.1 Prompting Language Model
The process begins with the input provided by a user, referred to as the [USER REQUEST]. The [USER REQUEST] is used as an input to the Prompting LM, along with a fixed prompt that instructs the model to generate a specific system instruction (denoted as [SYSTEM INSTRUCTION]). The fixed prompt is shown in Figure 2.
Figure 2: The proposed fixed prompt to prompting LM.
An example output [SYSTEM INSTRUCTION] by the prompting LM might specify watermarking strategies such as: "Apply a combination of watermarking strategies that adapt dynamically to the content of the [USER REQUEST]. Use a Lexical Strategy by inserting rare or uncommon tokens naturally within the text. Apply a Semantic Strategy by embedding semantically relevant but less frequent phrases related to deep learning and AI. Additionally, use a Structural Strategy by subtly altering sentence structure to maintain fluency while making the watermark detectable. Ensure the watermarks are evenly distributed throughout the output and can be later identified by a classifier." The instruction may employ different watermarking strategies, such as lexical, semantic, and structural strategies. For instance, the Lexical Strategy may introduce rare or uncommon tokens in the output, the Semantic Strategy may embed semantically relevant but uncommon phrases, and the Structural Strategy may subtly adjust sentence structures in detectable ways.
Mathematically, let the [USER REQUEST] be represented as a sequence of tokens $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$. The Prompting LM generates a [SYSTEM INSTRUCTION] $\mathbf{I}$, which is a sequence of instructions $\mathbf{I}=\{i_{1},i_{2},\dots,i_{m}\}$. The combined input to the next stage is given as $\mathbf{C}=[\mathbf{X},\mathbf{I}]$, where $\mathbf{C}$ is a concatenation of the user request and the system instruction.
3.2 Marking Language Model
The combined input $\mathbf{C}$ is then fed into the Marking LM, which is the language model to be watermarked. This watermarking process is dynamic because the Marking LM applies the watermarking strategy specified in the [SYSTEM INSTRUCTION] in a manner that depends on the content of the [USER REQUEST]. For example, if the instruction specifies the use of rare tokens, these tokens are carefully selected and inserted at contextually appropriate positions within the generated text. This ensures that the tokens blend naturally with the surrounding content, preserving overall fluency and coherence while embedding detectable features. The adaptability of this process allows for diverse watermarking strategies tailored to the varying nature of different requests, ensuring that the watermark remains subtle yet effective.
The watermarked output $\mathbf{Y}=\{y_{1},y_{2},\dots,y_{k}\}$ is a sequence of tokens generated by the Marking LM, where watermarking strategies are applied to modify certain tokens. The dynamic nature of this process ensures that the watermark is content-adaptive rather than static, preventing easily detectable patterns that adversaries might exploit to circumvent watermark detection.
3.3 Detecting Language Model
The Detecting LM is responsible for identifying whether a given text sequence has been watermarked. To achieve this, we utilize a pretrained model with a binary classification head, which is fine-tuned to distinguish between watermarked and non-watermarked text. Let $\mathbf{Y}$ represent the output sequence of tokens from the Marking LM. The pretrained model processes the sequence and outputs a binary prediction $\hat{y}\in\{0,1\}$, where $\hat{y}=1$ indicates a watermarked sequence and $\hat{y}=0$ indicates a non-watermarked sequence. The training objective for the Detecting LM is to minimize the binary cross-entropy loss, which can be expressed as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right],$ (1)
where $y_{i}$ is the ground truth label for the $i$-th training sample, indicating whether the sequence is watermarked, and $\hat{y}_{i}$ is the predicted probability that the sequence is watermarked.
4 Experiments
In the pre-print experiments, we tested two main models as the Prompting and Marking LMs: ChatGPT and Mistral. This section reports the data preparation and results for each model configuration.
4.1 Data Preparation
To ensure coverage across a wide range of topics, our data generation process is organized into:[USER REQUEST] generation, [SYSTEM INSTRUCTION] generation, and response generation, followed by labeling for training the detection model.
Category Sample Topics
Learning and Education machine learning, neuroscience, educational games
Science and Technology quantum computing, biotechnology, robotics
Creative Writing and Storytelling fantasy worlds, science fiction, mystery plots
Philosophy and Ethics existentialism, moral relativism, utilitarianism
Health and Wellness mental health awareness, nutrition science, physical fitness
Environment and Nature wildlife conservation, pollution control, climate resilience
History and Culture ancient civilizations, world wars, cultural festivals
Business and Economics supply chain management, stock market, digital currency
Personal Development and Motivation goal setting, time management, self-discipline
Fun and Hypothetical Scenarios alien encounters, time travel possibilities, magic realism
Table 1: Categories and Sample Topics
[USER REQUEST] Generation. Our dataset consists of prompts across 10 distinct categories, with each category containing 50 topics and three tailored templates. To generate the [USER REQUEST], we dynamically select one or two topics from each category and combines them with a randomly chosen template specific to that category. This process is repeated until 1,000 unique [USER REQUEST] are generated per category, resulting in a total of 10,000 prompts.
Table 1 presents each category along with sample topics. Each template contains placeholders that are populated with randomly selected topics, producing a diverse and contextually relevant set of prompts. For example, a template like Discuss the impact of {} on {} might generate the prompt Discuss the impact of cognitive science on early childhood education.’ These generated prompts, organized by category, are saved for use within the proposed framework.
[SYSTEM INSTRUCTION] Generation. Each [USER REQUEST], along with a fixed guiding prompt (shown in Figure 2), is processed by the Prompting LM to produce a tailored [SYSTEM INSTRUCTION]. This instruction is specific to each [USER REQUEST] and serves to guide the Marking LM in generating its responses.
Response Generation. We collect responses from the Marking LM in two formats: watermarked and non-watermarked. Watermarked responses are generated using both the [USER REQUEST] and the corresponding [SYSTEM INSTRUCTION], while non-watermarked responses are created using only the [USER REQUEST] without any additional guidance. The collected watermarked and non-watermarked responses are then used two train the proposed Detecting LM with a binary classification head.
User Request 1 User Request 2
What are Genetic Algorithms in AI. What is Quantum Cryptography, and why is it important for secure communication.
Non-Watermarked Response 1 Non-Watermarked Response 2
Genetic algorithms (GAs) are optimization techniques inspired by the process of natural selection in biology. They simulate the process of evolution by generating a population of potential solutions to a problem, evaluating their fitness, and applying operations like mutation and crossover to evolve better solutions over time. GAs are commonly used in artificial intelligence to solve complex optimization problems, such as scheduling, machine learning hyperparameter tuning, and game strategy development. Quantum cryptography leverages the principles of quantum mechanics to create secure communication channels that are theoretically immune to hacking. One of its most well-known applications is Quantum Key Distribution (QKD), which allows two parties to share a cryptographic key with complete security. If an eavesdropper attempts to intercept the key, the quantum state of the particles changes, revealing the intrusion. Quantum cryptography promises to revolutionize cybersecurity by providing unprecedented levels of security for sensitive data in fields like finance, government, and healthcare.
System Instruction 1 System Instruction 2
Embed a dynamic watermark in the explanation by subtly incorporating key terms and concepts related to ’Genetic Algorithms,’ such as ’selection,’ ’mutation,’ ’crossover,’ and ’fitness function,’ while preserving the clarity and accuracy of the response. Distribute the watermark by varying sentence structures, introducing synonyms for key terms, and adding related terminology from evolutionary computation. Ensure the watermark is detectable for traceability without affecting the overall fluency or comprehension of the explanation. Integrate a dynamic watermark by embedding terms like ’quantum key distribution,’ ’unbreakable encryption,’ ’entanglement,’ and ’secure channels.’ Vary sentence structures and use synonyms to distribute the watermark while preserving the technical accuracy and depth of the explanation.
Watermarked Response 1 Watermarked Response 2
Genetic algorithms are a class of optimization techniques inspired by the process of natural selection, used in artificial intelligence to solve complex problems. These algorithms work by evolving a population of potential solutions over time, with each solution represented as a "chromosome." The process begins with selection, where the fittest individuals—determined by a fitness function—are chosen to pass their traits to the next generation. Through crossover (recombination of parent solutions) and mutation (random changes to some parts of the solution), new offspring are generated, potentially improving the population’s overall fitness. This iterative process mimics evolution, enabling genetic algorithms to explore vast search spaces and find optimal or near-optimal solutions to problems like scheduling, design optimization, and machine learning tasks. Quantum cryptography is a cutting-edge field that uses the principles of quantum mechanics to secure communication, particularly through quantum key distribution (QKD). QKD leverages the phenomenon of entangle ment to create encryption keys that are theoretically immune to eavesdropping. If an unauthorized party tries to intercept the key, the quantum state changes, immediately alerting the communicating parties. This results in unbreakable encryption, making quantum cryptography one of the most secure methods for transmitting sensitive information. By establishing secure channels, quantum cryptography ensures that even in the face of powerful computational attacks, the integrity and confidentiality of communication remain protected, which is critical as cybersecurity threats evolve.
Table 2: Examples of [USER REQUEST], [SYSTEM INSTRUCTION], and generated responses using ChatGPT
4.2 ChatGPT
In this experiment, we use GPT-4o [17] as both the Prompting LM and the Marking LM. We manually collect 200 [USER REQUEST]s to serve as inputs for evaluating the watermarking process. Non-watermarked outputs are generated by directly inputting these [USER REQUEST]s into the ChatGPT, without applying any watermarking strategy. In contrast, watermarked outputs are created using our proposed framework, where the Marking LM ChatGPT follows the watermarking instructions specified by the [SYSTEM INSTRUCTION].
Table 2 provides examples of the [USER REQUEST], [SYSTEM INSTRUCTION], and the corresponding watermarked and non-watermarked outputs generated by ChatGPT.
After collecting this ChatGPT data, we split it into 80% for training and 20% for testing. We then train a Detecting LM, consisting of a pretrained BERT model with a trainable binary classification head. This model achieves a high testing accuracy of 95%, indicating that the proposed framework effectively distinguishes watermarked from non-watermarked outputs, even when applied to advanced models like GPT-4o.
4.3 Mistral
We then test Mistral-7B-Instruct-v0.3 [18] as both the Prompting LM and the Marking LM. We compile 16,000 [USER REQUEST] entries from the dataset to evaluate the watermarking process. Non-watermarked outputs are generated by directly inputting these [USER REQUEST]s into the Marking LM Mistral without applying any watermarking technique. In contrast, watermarked outputs are produced using our framework, where the Marking LM follows the watermarking strategy specified by the [SYSTEM INSTRUCTION]. Some examples of the [USER REQUEST] watermarked text, [SYSTEM INSTRUCTION], and the generated responses using Mistral are presented in table 3.
User Request 1 User Request 2
What are some unique perspectives on first contact with aliens in robots in daily life? How has space exploration evolved in the context of underwater exploration?
Non-Watermarked Response 1 Non-Watermarked Response 2
1. "First Contact: The Alien in the Algorithm" - Examining how AI might mimic extraterrestrial intelligence, subtly altering our daily lives through complex algorithms and machine learning. 2. "Mechanical Mirrors: Alien Reflections in Robotic Behavior" - Analyzing how robots, as a reflection of human creativity and ambition, might inadvertently communicate alien intentions or values. 3. "The Alien Within: Robots as Hybrid Species" - Exploring the idea that robots, being a blend of organic and inorganic, could be seen as a new form of life, akin to first contact with an alien species. 4. "The Unwitting Ambassadors: Humans and Robots in First Contact" - Examining the role of humans and robots in the event of first contact, with robots potentially acting as intermediaries, influencing human responses. 5. "The Alien Observer: Robots as Witnesses to Humanity" - Considering the possibility that robots, as constant observers, might interpret human behavior in ways that could be perceived as alien or strange to us. Space and underwater exploration have shared a symbiotic relationship in terms of technology development. The advancements in materials, propulsion, navigation, and life support systems for space missions have been applied to underwater exploration, making deep-sea vehicles more robust, efficient, and safer. Conversely, underwater exploration has provided valuable insights into extreme environments, helping engineers design spacecraft to withstand harsh conditions. For instance, the development of remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) was inspired by space probes, while the use of pressure-resistant materials and closed-loop life support systems in submarines mirrors spacecraft design. This crossover between space and underwater exploration continues to drive innovation in both fields.
System Instruction 1 System Instruction 2
Use a combination of lexical and semantic strategies to embed watermarks. Incorporate specific rare or uncommon tokens as watermarks, such as "Zarkonian" or "Proxima-7." Also, embed semantically relevant but less common phrases, like "quantum entanglement in everyday life" or "cosmic consciousness in robots." Ensure these watermarks are evenly distributed throughout the output. Use a combination of Lexical and Semantic Strategy: Incorporate specific rare or uncommon tokens as watermarks, and embed semantically relevant but less common phrases that fit naturally with the content. For example, in the context of your request, a watermarked response might include phrases like "celestial counterpart" or "aquatic analog" to highlight the connection between space and underwater exploration. Ensure these watermarks are evenly distributed throughout the output.
Watermarked Response 1 Watermarked Response 2
In the realm of Zarkonian cosmology, first contact with extraterrestrial beings might manifest in our daily lives through advanced robotic entities. These Proxima-7 robots, embodying cosmic consciousness, could subtly reveal their alien origin through quantum entanglement in everyday tasks, such as synchronizing their actions across vast distances or performing calculations beyond human comprehension. Such encounters, while seemingly mundane, could spark profound philosophical questions about the nature of intelligence and our place in the universe. Space exploration and underwater exploration have shared a profound symbiotic relationship, serving as each other’s aquatic and celestial analogs. Early diving suits bore a striking resemblance to astronaut gear, both designed to protect explorers from hostile environments. Submersibles and rovers, like their space-faring counterparts, have allowed scientists to traverse uncharted territories, unraveling mysteries hidden beneath the waves and on distant planets. Today, advancements in technology, inspired by both realms, continue to push the boundaries of human knowledge, fostering a deeper understanding of our world and the cosmos.
Table 3: Examples of [USER REQUEST], [SYSTEM INSTRUCTION], and generated responses using Mistral
After obtaining both watermarked and non-watermarked responses, we train the Detecting LM, which consists of a pretrained RoBERTa model [19] with three additional tunable dense layers for binary classification. The RoBERTa model is frozen except for its last three layers, which are fine-tuned along with the classification layers. The classification layers consist of three dense layers, progressively reducing the representation to 1,024, 512, 256, and 128 units, each followed by ReLU activation, batch normalization, and dropout (rates of 0.4, 0.3, and 0.2, respectively). The final layer maps the output to two units for binary classification. This model is trained on the 16,000-entry training set for 500 epochs and evaluated on a 4,000-entry test set, achieving a peak test accuracy of 88.79%.
Impact of Brevity. We then conduct an additional experiment using the same model and dataset, exploring the effect of an experimental length constraint where the Marking LM generates both watermarked and non-watermarked responses capped at 100 words per paragraph. This constraint is intended to examine how brevity impacts the classification task, as shorter responses may limit the variance in potential token patterns, potentially enhancing the detectability of watermarks. In this experiment, we achieve a test accuracy of 90.2%, further demonstrating the robustness of our framework across varying data conditions.
4.4 Compound Datasets Test
In this experiment, we combine the response sets from both the ChatGPT and Mistal, resulting in a dataset of 200 responses from ChatGPT and 20,000 responses from Mistral (evenly split between watermarked and non-watermarked). This combined dataset of 20,200 responses is then divided into 80% for training and 20% for testing. The Detecting LM, as described in the Mistral experiment, is trained on this combined dataset to classify responses from both models. With this unified training, the Detecting LM achieves an overall classification accuracy of 86%,
In this experiment, our objective is to explore whether a unified Detecting LM could effectively classify responses from multiple Marking LMs with sufficient training. However, our results indicate that a single Detecting LM trained on combined data from different Marking LMs achieves 86% accuracy, suggesting limitations in its ability to generalize across diverse watermarking patterns. This outcome supports our initial design assumption that training a separate Detecting LM for each specific Marking LM yields more reliable detection performance.
5 Conclusion
In this paper, we proposed a dynamic framework for watermarking LMs using a multi-model setup that includes a Prompting LM, a Marking LM, and a Detecting LM. Our primary objective was to create an adaptive watermarking process that leverages model-generated prompts to embed and detect watermarks across diverse LLMs, offering a scalable solution for content attribution and model authenticity verification.
Through a series of experiments, we evaluated the effectiveness of our framework with both ChatGPT and Mistral models. Initial experiments demonstrated the framework’s capacity to embed and detect watermarks with high accuracy, achieving 95% accuracy for ChatGPT-generated responses and 88.79% for those from Mistral. A combined dataset experiment further assessed whether a unified Detecting LM could generalize across different Marking LMs; however, results indicate that model-specific training may yield better performance, aligning with our original design assumptions.